OpenStudy (anonymous):

Given a circle with a radius of 6, what is the length of an arc measuring 60°?

OpenStudy (anonymous):

i got 44/7

OpenStudy (binary3i):

arc length= angle* ragius =60*6 =pi*6/3=2pi

OpenStudy (phi):

2 pi If you use pi is approximate ly 22/7 then 44/7

OpenStudy (binary3i):

oh yeah.